# Compositional Reasoning with Transformers, RNNs, and Chain of Thought

### v1

@@ Appendix B Memory-rank sort
In this section we formally define the memory-rank sort. It is presented in Algorithm 1. We next prove that performing memory-rank sort and calculting the memory-rank of all the nodes in a tree can be done efficiently:
@@
The time that Algorithm 1 runs is the number of iterations in the “while” loop. Note that each leaf in the tree is visited at most once, since if $v_{\text{cur}}$ is a leaf, it is inserted to the list and not visited again. For any non-leaf node, it is visited at most $3$ times during the loop. Once for each one of its children, and again when it is inserted into the list. In total the running time of the sorting algorithm is bounded by $3|V|$. ∎
@@
Appendix C Proofs from Section 5
C.1 Proof of Theorem 5.2
@@ Appendix C Proofs from Section 5
For simplicity of the proof, we use trees with $14n-$ 1 nodes instead of $n$, this will only effect the constant in the $\Omega$ notation. We construct the following CRQ: The tree $T=(V,E)$ is a balanced binary tree with $14n-$ 1 nodes, thus it has $2n$ leaves. All the values of the leaves are in {±$\{\pm 1\}$. The nodes in the second to last layer are all equal to $-1$, and the rest non-leaf nodes are equal to $1$. We number the leaves as $v_{1},\dots,v_{2n}$. Suppose we are given an instance of the set-disjointness problem, where Alice is given the leaves $v_{2i-1}$ and Bob the leaves $v_{2i}$ for $i\in[n]$. By the definition of CRQ, the answer to the second to last layer nodes with leaves $v_{2i-1}$ and $v_{2i}$ is $1$ if $v_{2i-1}=v_{2i}=1$ and $-1$ otherwise. In addition the answer to any other node will be $1$ if one of its children is $1$ and $-1$ otherwise. In total, the answer to the tree is $1$ if there exists some $i\in[n]$ with $v_{2i-1}=v_{2i}=1$ and $-1$ otherwise.
Consider the following ordering of the nodes for the RNN: The first $n$ nodes are $v_{2i-1}$ for $i\in[n]$, the next $n$ nodes are $v_{2i}$ for $i\in[n]$. The root is the last node in this ordering, and the rest of the nodes are given in some arbitrary order. Suppose there is an RNN known to both Alice and Bob that solves the CRQ problem above in the prescribed order of the nodes, and it has a hidden state $\mathbf{h}\in\mathbb{R}^{m}$ where each coordinate is represented by $p$ bits. We will define the following communication protocol to solve the set disjointness problem: Alice apply the RNN on her inputs. After each input $v_{2i-1}$ for $i\in[n]$ she passes the hidden state $\mathbf{h}_{i}$ to the next recurrence of the RNN. After $n$ such recurrences of the RNN, the hidden state $\mathbf{h}_{n}$ is passed to Bob, he inputs his inputs $v_{2j}$ for $j\in[n]$ in the prescribed order, each time passing a hidden state. He continue to run the RNN on the rest of the nodes, and output the answer to the CRQ. The number of bits transferred between Alice and Bob is $m\cdot p$, since the only communication between them is transferring $\mathbf{h}_{n}$ and Alice finishes processing her inputs. By C.1 we have that $mp=\Omega(n)$, which finishes the proof.
C.2 Proof of Theorem 5.4
@@
Let $r_{k}$ be the smallest number of descendants for a node with memory-rank $k$. It is easy to see that $r_{1}=2$, which happens when the node has only two children that are leaves. Let $v$ be a node with $\textbf{mr}(v)=k$, and let $u_{1},u_{2}$ be its children. It can be seen that $r_{k}\geq 2r_{k-1}$. This is because, if $\textbf{mr}(u_{1})\neq\textbf{mr}(u_{2})$, then one of them must be equal to $k$, assume w.l.o.g it is $u_{1}$. This means that either $u_{1}$ or one of its descendants have two children with a memory-rank of $k-1$. In particular, the number of descendants of $v$ is larger than $2r_{k-1}$. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, then they are both equal to $k-1$, in which case the same conclusion follows. Applying the recurrence formula, we get that $r_{k}\geq 2^{k}$.
@@
Let $R,\epsilon>0$ and $d\in\mathbb{N}$. There exists a $2$-layer neural network $\mathcal{N}:\mathbb{R}^{2d}\rightarrow\mathbb{R}$ with width $O\left(\frac{d^{2}R}{\epsilon}\right)$ such that $\max_{\mathbf{x},\mathbf{y}\in[-R,R]^{d}}|\mathcal{N}(\mathbf{x},\mathbf{y})-\left\langle\mathbf{x},\mathbf{y}\right\rangle|\leq\epsilon$
@@
We now turn to the formal construction. Each node $v\in V$ in the tree will be embedded as a vector $\begin{pmatrix}\mathbf{x}_{v}\\ \ell(v)\end{pmatrix}\in\mathbb{R}^{d+1}$, where $\ell(v)$ is the depth of $v$. The hidden state of the RNN at the beginning will be $\mathbf{h}=\bm{0}_{d(\log(n)+1)}$. The input to the RNN will be the current input vector, concatenated to the hidden state. Throughout the proof we refer to $\tilde{\mathbf{x}}$ the vector that is inputted to the RNN, namely the concatenation of the embedding of each node with the hidden state. We call the part of $\tilde{\mathbf{x}}$ that contains the hidden state $\mathbf{h}$ the stack, and each consecutive $d+1$ coordinates in it as a node that is was inputted to the stack. The reason is that this part of the input will simulate a stack where the RNN can only access its first and second nodes, namely the first $22d+$ 2 coordinates. The RNN will add at most $3$ coordinates to $\tilde{\mathbf{x}}$, which will simulate flags with values that are either 00 or $1$. The RNN contains the following layers:
@@
First layer: We denote the input vector to the RNN is $\tilde{\mathbf{x}}$. The first layer add a coordinate to the input. All the input coordinates are copied, while the last coordinate is constructed as: $\sigma((\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}+1)-\sigma((\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1})$. Note that this added last coordinate is equal to $1$ if $(\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}\geq 0$, meaning that the current node has an equal or larger depth that the last node in the stack, and 00 if $(\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}\leq-1$. The copying of the input can be done by a ReLU network, since $\sigma(z)-\sigma(-z)=z$ for every $z\in\mathbb{R}$. Thus, adding two identity matrices to the weights of the MLP with different signs, and adding their outputs together will copy the inputs.
@@
This operation can be implemented by a $2$-layer network, where the constant $\boldsymbol{1}$ vectors are added using the bias terms. This operation will apply the identity matrix to the inputs if the flag (defined in the previous layer) is 00, and apply the $A$ matrix if it is $1$. The $A$ matrix will copy the first $d+1$ entries (the current input node), and will move also add it to the stack, while also moving all the other entries $d+1$ coordinates further down in the stack. The added $c\boldsymbol{1}$ factor is so that if there are negative values in the entries of the inputs, they will not be removed by the ReLU, this factor is removed at the end to keep the original value. We will later on prove that the stack will not get overflown, meaning that we don’t delete the last $d+1$ entries of it.
@@
By Lemma C.3 there exists a $2$-layer neural network $\mathcal{N}:\mathbb{R}^{2d}\rightarrow\mathbb{R}$ with width $O(d^{2})$ such that $\max_{\mathbf{x},\mathbf{y}\in\Gamma^{d}}|\mathcal{N}(\mathbf{x},\mathbf{y})-\left\langle\mathbf{x},\mathbf{y}\right\rangle|\leq\frac{\epsilon}{4}$. We will stack two such networks that approximate up to an error $\frac{\epsilon}{4}$ the inner products $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle,\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle$. If $\mathbf{x}_{u_{1}}\neq\mathbf{x}_{u_{2}}$, then by the assumption on $\epsilon$ we have that $|\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle-\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle|\geq\epsilon$. By the construction of $\mathcal{N}$, we also have that:
@@
Fourth and Fifth layers: In the last layer, we add an additional flag on whether $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle>\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle$ and 00 otherwise. This can be calculated by using the following function:
@@
By the argument from the previous layer, this function will provide the additional flag. All the other coordinates are copied as is. We use another layer to pop out the first $22d+$ 2 coordinates from the stack, meaning that all the entries in the stack are moved upward by $d+1$ coordinates, and the first $d+1$ are changed to either $\begin{pmatrix}\mathbf{x}_{u_{1}}\\ \ell(u_{1})-1\end{pmatrix}$ or $\begin{pmatrix}\mathbf{x}_{u_{2}}\\ \ell(u_{2})-1\end{pmatrix}$, depending on whether the new flag is $1$ or 00. Again, we apply this operation only if the flag from the first layer is turned off, by using a similar construction to the second layer. The rest of the coordinates are copied, and the coordinates of the flags that were used throughout the computation are removed. The output will be the hidden state to the next recurrence with the dimension as the hidden state that was inputted.
@@
We are left with showing that the stack does not overflow. We will show that the maximal number of nodes in the stack is bounded by $\textbf{mr}(T)+1$. Since the number of coordinates in the representation of each node is $d+1$, using Lemma C.2 finishes the proof. Let $v$ be some node with $\textbf{mr}(v)=k$. We will first show by induction on $k$ that at any point in time, before processing $v$, the maximal number of nodes in the stacks that are also descendants of $v$ is bounded by $k$. If $k=0$ then $v$ is a leaf a doesn’t have any children. For the induction step, assume $k>0$, and let $u_{1},u_{2}$ be the two children of $v$. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, then by the definition of memory rank, they are both equal to $k-1$. Assume w.l.o.g that $u_{1}$ is processed before $u_{2}$. Note that in this case, $u_{2}$ is processed only after all the descendants of $u_{1}$ are processed. Then, by the induction step, the maximal number of nodes in the stacks that are also descendants of $u_{2}$ is bounded by $k-1$. Hence, the maximal number of nodes in the stack that are descendants of $v$ is $k$, which includes $u_{1}$. If $\textbf{mr}(u_{1})\neq\textbf{mr}(u_{2})$, assume w.l.o.g that $\textbf{mr}(u_{1})>\textbf{mr}(u_{2})$, then $u_{1}$ is processed before $u_{2}$ (resp. $u_{2}$ before $u_{1}$). In this case, by the induction case, again the number of nodes in the stack the are descendants of $v$, are either descendants of $u_{1}$, in which case no nodes that are descendants of $u_{2}$ have being processed, or descendants of $u_{2}$, in which case $u_{1}$ is in the stack, as well as at most $k-1$ descendants of $u_{2}$. This finished the induction.
@@
By induction on $\textbf{mr}(v)$. If $\textbf{mr}(v)=0$, then it is a leaf. By Algorithm 1, since $v$ doesn’t have children, the only node that is processed before it can be its sibling. When its sibling was processed necessarily all its children were already processed, hence there is at most $1$ node in the stack. Assume that if $\textbf{mr}(u)=k$, then there are at most $k+1$ nodes in the stack, and let $v$ with $\textbf{mr}(v)=k+1$. Denote by $u_{1}$ and $u_{2}$ the children of $v$, and assume w.l.o.g that $\textbf{mr}(u_{1})\geq\textbf{mr}(u_{2})$. If $\textbf{mr}(u_{1})>\textbf{mr}(u_{2})$, then $\textbf{mr}(v)=\textbf{mr}(u_{1})$. This means that $u_{1}$ was processed before $u_{2}$. Hence, $u_{1}$ is in the stack, and $u_{2}$ was processed right before $v$, which by induction shows that there are at most $k+1$ nodes in the stack. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, the by the definition of memory-rank, it necessarily happen that they are both equal to $k$. Assume w.l.o.g that $u_{1}$ is processed before $u_{2}$ (the order of them being processed is chosen arbitrarily in this case). Hence, when processing $u_{2}$ there are at most $k$ nodes. ∎
C.3 Proof of Theorem 5.5
@@
As in Lemma 4.2, an intuitive way to look at these CRQs is as Boolean formulas. For leaf nodes $\mathbf{t}_{0}$ represents 00 and $\mathbf{t}_{1}$ represents $1$. For non-leaf nodes, $\mathbf{t}_{0}$ represents $\land$ and $\mathbf{t}_{1}$ represents $\lor$. The vector $\mathbf{t}_{\texttt{null}}$ is used by a hidden node to pass values up the tree without changing them, as described below.
@@
Now, let $T$ be some rooted tree of size $n$, which is non necessarily a balanced binary tree. Denote by $\ell:=\textbf{mr}(T)$, and recall this means that the root has a memory-rank of $\ell$. We will find a subset of $T$ that forms a balanced binary tree whose depth equals $\ell$. We will construct all the other nodes to simply copy their inputs up the tree, and appeal to the case of a balanced binary tree proved above.

### v2

@@ Appendix B Memory-rank sort
In this section we formally define the memory-rank sort. It is presented in Algorithm˜1. We next prove that performing memory-rank sort and calculting the memory-rank of all the nodes in a tree can be done efficiently:
@@
The time that Algorithm˜1 runs is the number of iterations in the “while” loop. Note that each leaf in the tree is visited at most once, since if $v_{\text{cur}}$ is a leaf, it is inserted to the list and not visited again. For any non-leaf node, it is visited at most $3$ times during the loop. Once for each one of its children, and again when it is inserted into the list. In total the running time of the sorting algorithm is bounded by $3|V|$. ∎
@@
Appendix C Proofs from Section˜5
C.1 Proof of Theorem˜5.2
@@ Appendix C Proofs from Section˜5
For simplicity of the proof, we use trees with 4n14n-1 nodes instead of $n$, this will only affect the constant in the $\Omega$ notation. We construct the following CRQ: The tree $T=(V,E)$ is a balanced binary tree with 4n14n-1 nodes, thus it has $2n$ leaves. All the values of the leaves are in $\{\pm 1\}$. The nodes in the second to last layer are all equal to 1-1, and the rest non-leaf nodes are equal to $1$. We number the leaves as $v_{1},\dots,v_{2n}$. Suppose we are given an instance of the set-disjointness problem, where Alice is given the leaves $v_{2i-1}$ and Bob the leaves $v_{2i}$ for $i\in[n]$. By the definition of CRQ, the answer to the second to last layer nodes with leaves $v_{2i-1}$ and $v_{2i}$ is $1$ if $v_{2i-1}=v_{2i}=1$ and 1-1 otherwise. In addition the answer to any other node will be $1$ if one of its children is $1$ and 1-1 otherwise. In total, the answer to the tree is $1$ if there exists some $i\in[n]$ with $v_{2i-1}=v_{2i}=1$ and 1-1 otherwise.
Consider the following ordering of the nodes for the RNN: The first $n$ nodes are $v_{2i-1}$ for $i\in[n]$, the next $n$ nodes are $v_{2i}$ for $i\in[n]$. The root is the last node in this ordering, and the rest of the nodes are given in some arbitrary order. Suppose there is an RNN known to both Alice and Bob that solves the CRQ problem above in the prescribed order of the nodes, and it has a hidden state $\mathbf{h}\in\mathbb{R}^{m}$ where each coordinate is represented by $p$ bits. We will define the following communication protocol to solve the set disjointness problem: Alice apply the RNN on her inputs. After each input $v_{2i-1}$ for $i\in[n]$ she passes the hidden state $\mathbf{h}_{i}$ to the next recurrence of the RNN. After $n$ such recurrences of the RNN, the hidden state $\mathbf{h}_{n}$ is passed to Bob, he inputs his inputs $v_{2j}$ for $j\in[n]$ in the prescribed order, each time passing a hidden state. He continue to run the RNN on the rest of the nodes, and output the answer to the CRQ. The number of bits transferred between Alice and Bob is $m\cdot p$, since the only communication between them is transferring $\mathbf{h}_{n}$ and Alice finishes processing her inputs. By ˜C.1 we have that $mp=\Omega(n)$, which finishes the proof.
C.2 Proof of Theorem˜5.4
@@
Let $r_{k}$ be the smallest number of descendants for a node with memory-rank $k$. It is easy to see that $r_{1}=2$, which happens when the node has only two children that are leaves. Let $v$ be a node with $\textbf{mr}(v)=k$, and let $u_{1},u_{2}$ be its children. It can be seen that $r_{k}\geq 2r_{k-1}$. This is because, if $\textbf{mr}(u_{1})\neq\textbf{mr}(u_{2})$, then one of them must be equal to $k$, assume w.l.o.g it is $u_{1}$. This means that either $u_{1}$ or one of its descendants have two children with a memory-rank of k1k-1. In particular, the number of descendants of $v$ is larger than $2r_{k-1}$. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, then they are both equal to k1k-1, in which case the same conclusion follows. Applying the recurrence formula, we get that $r_{k}\geq 2^{k}$.
@@
Let $R,\epsilon>0$ and $d\in\mathbb{N}$. There exists a $2$-layer neural network $\mathcal{N}:\mathbb{R}^{2d}\rightarrow\mathbb{R}$ with width $O\left(\frac{d^{2}R}{\epsilon}\right)$ such that $\max_{\mathbf{x},\mathbf{y}\in[-R,R]^{d}}|\mathcal{N}(\mathbf{x},\mathbf{y})-\left\langle\mathbf{x},\mathbf{y}\right\rangle|\leq\epsilon$ .
@@
We now turn to the formal construction. Each node $v\in V$ in the tree will be embedded as a vector $\begin{pmatrix}\mathbf{x}_{v}\\ \ell(v)\end{pmatrix}\in\mathbb{R}^{d+1}$, where $\ell(v)$ is the depth of $v$. The hidden state of the RNN at the beginning will be $\mathbf{h}=\bm{0}_{d(\log(n)+1)}$. The input to the RNN will be the current input vector, concatenated to the hidden state. Throughout the proof we refer to $\tilde{\mathbf{x}}$ the vector that is inputted to the RNN, namely the concatenation of the embedding of each node with the hidden state. We call the part of $\tilde{\mathbf{x}}$ that contains the hidden state $\mathbf{h}$ the stack, and each consecutive $d+1$ coordinates in it as a node that is was inputted to the stack. The reason is that this part of the input will simulate a stack where the RNN can only access its first and second nodes, namely the first $2d+2$ coordinates. The RNN will add at most $3$ coordinates to $\tilde{\mathbf{x}}$, which will simulate flags with values that are either 0 or $1$. The RNN contains the following layers:
@@
First layer: We denote the input vector to the RNN is $\tilde{\mathbf{x}}$. The first layer add a coordinate to the input. All the input coordinates are copied, while the last coordinate is constructed as: $\sigma((\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}+1)-\sigma((\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1})$. Note that this added last coordinate is equal to $1$ if $(\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}\geq 0$, meaning that the current node has an equal or larger depth that the last node in the stack, and 0 if $(\tilde{\mathbf{x}})_{d+1}-(\tilde{\mathbf{x}})_{2d+1}\leq-1$. The copying of the input can be done by a ReLU network, since $\sigma(z)-\sigma(-z)=z$ for every $z\in\mathbb{R}$. Thus, adding two identity matrices to the weights of the MLP with different signs, and adding their outputs together will copy the inputs.
@@
This operation can be implemented by a $2$-layer network, where the constant $\boldsymbol{1}$ vectors are added using the bias terms. This operation will apply the identity matrix to the inputs if the flag (defined in the previous layer) is 0, and apply the $A$ matrix if it is $1$. The $A$ matrix will copy the first $d+1$ entries (the current input node), and will move also add it to the stack, while also moving all the other entries $d+1$ coordinates further down in the stack. The added $c\boldsymbol{1}$ factor is so that if there are negative values in the entries of the inputs, they will not be removed by the ReLU, this factor is removed at the end to keep the original value. We will later on prove that the stack will not get overflown, meaning that we don’t delete the last $d+1$ entries of it.
@@
By Lemma˜C.3 there exists a $2$-layer neural network $\mathcal{N}:\mathbb{R}^{2d}\rightarrow\mathbb{R}$ with width $O(d^{2})$ such that $\max_{\mathbf{x},\mathbf{y}\in\Gamma^{d}}|\mathcal{N}(\mathbf{x},\mathbf{y})-\left\langle\mathbf{x},\mathbf{y}\right\rangle|\leq\frac{\epsilon}{4}$. We will stack two such networks that approximate up to an error $\frac{\epsilon}{4}$ the inner products $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle,\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle$. If $\mathbf{x}_{u_{1}}\neq\mathbf{x}_{u_{2}}$, then by the assumption on $\epsilon$ we have that $|\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle-\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle|\geq\epsilon$. By the construction of $\mathcal{N}$, we also have that:
@@
Fourth and Fifth layers: In the last layer, we add an additional flag on whether $\left\langle\mathbf{x}_{u_{1}},\mathbf{x}_{v}\right\rangle>\left\langle\mathbf{x}_{u_{2}},\mathbf{x}_{v}\right\rangle$ and 0 otherwise. This can be calculated by using the following function:
@@
By the argument from the previous layer, this function will provide the additional flag. All the other coordinates are copied as is. We use another layer to pop out the first $2d+2$ coordinates from the stack, meaning that all the entries in the stack are moved upward by $d+1$ coordinates, and the first $d+1$ are changed to either $\begin{pmatrix}\mathbf{x}_{u_{1}}\\ \ell(u_{1})-1\end{pmatrix}$ or $\begin{pmatrix}\mathbf{x}_{u_{2}}\\ \ell(u_{2})-1\end{pmatrix}$, depending on whether the new flag is $1$ or 0. Again, we apply this operation only if the flag from the first layer is turned off, by using a similar construction to the second layer. The rest of the coordinates are copied, and the coordinates of the flags that were used throughout the computation are removed. The output will be the hidden state to the next recurrence with the dimension as the hidden state that was inputted.
@@
We are left with showing that the stack does not overflow. We will show that the maximal number of nodes in the stack is bounded by $\textbf{mr}(T)+1$. Since the number of coordinates in the representation of each node is $d+1$, using Lemma˜C.2 finishes the proof. Let $v$ be some node with $\textbf{mr}(v)=k$. We will first show by induction on $k$ that at any point in time, before processing $v$, the maximal number of nodes in the stacks that are also descendants of $v$ is bounded by $k$. If $k=0$ then $v$ is a leaf a doesn’t have any children. For the induction step, assume $k>0$, and let $u_{1},u_{2}$ be the two children of $v$. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, then by the definition of memory rank, they are both equal to k1k-1. Assume w.l.o.g that $u_{1}$ is processed before $u_{2}$. Note that in this case, $u_{2}$ is processed only after all the descendants of $u_{1}$ are processed. Then, by the induction step, the maximal number of nodes in the stacks that are also descendants of $u_{2}$ is bounded by k1k-1. Hence, the maximal number of nodes in the stack that are descendants of $v$ is $k$, which includes $u_{1}$. If $\textbf{mr}(u_{1})\neq\textbf{mr}(u_{2})$, assume w.l.o.g that $\textbf{mr}(u_{1})>\textbf{mr}(u_{2})$, then $u_{1}$ is processed before $u_{2}$ (resp. $u_{2}$ before $u_{1}$). In this case, by the induction case, again the number of nodes in the stack the are descendants of $v$, are either descendants of $u_{1}$, in which case no nodes that are descendants of $u_{2}$ have being processed, or descendants of $u_{2}$, in which case $u_{1}$ is in the stack, as well as at most k1k-1 descendants of $u_{2}$. This finished the induction.
@@
By induction on $\textbf{mr}(v)$. If $\textbf{mr}(v)=0$, then it is a leaf. By Algorithm˜1, since $v$ doesn’t have children, the only node that is processed before it can be its sibling. When its sibling was processed necessarily all its children were already processed, hence there is at most $1$ node in the stack. Assume that if $\textbf{mr}(u)=k$, then there are at most $k+1$ nodes in the stack, and let $v$ with $\textbf{mr}(v)=k+1$. Denote by $u_{1}$ and $u_{2}$ the children of $v$, and assume w.l.o.g that $\textbf{mr}(u_{1})\geq\textbf{mr}(u_{2})$. If $\textbf{mr}(u_{1})>\textbf{mr}(u_{2})$, then $\textbf{mr}(v)=\textbf{mr}(u_{1})$. This means that $u_{1}$ was processed before $u_{2}$. Hence, $u_{1}$ is in the stack, and $u_{2}$ was processed right before $v$, which by induction shows that there are at most $k+1$ nodes in the stack. If $\textbf{mr}(u_{1})=\textbf{mr}(u_{2})$, the by the definition of memory-rank, it necessarily happen that they are both equal to $k$. Assume w.l.o.g that $u_{1}$ is processed before $u_{2}$ (the order of them being processed is chosen arbitrarily in this case). Hence, when processing $u_{2}$ there are at most $k$ nodes. ∎
C.3 Proof of Theorem˜5.5
@@
As in Lemma˜4.2, an intuitive way to look at these CRQs is as Boolean formulas. For leaf nodes $\mathbf{t}_{0}$ represents 0 and $\mathbf{t}_{1}$ represents $1$. For non-leaf nodes, $\mathbf{t}_{0}$ represents $\land$ and $\mathbf{t}_{1}$ represents $\lor$. The vector $\mathbf{t}_{\texttt{null}}$ is used by a hidden node to pass values up the tree without changing them, as described below.
@@
Now, let $T$ be some rooted tree of size $n$, which is not necessarily a balanced binary tree. Denote by $\ell:=\textbf{mr}(T)$, and recall this means that the root has a memory-rank of $\ell$. We will find a subset of $T$ that forms a balanced binary tree whose depth equals $\ell$. We will construct all the other nodes to simply copy their inputs up the tree, and appeal to the case of a balanced binary tree proved above.